# GenExam: A Multidisciplinary Text-to-Image Exam

### v2

@@ Appendix C Additional Experiments
BLIP3o-NEXT-GRPO-Text-3B
Biology/Ecology 47.1 36.3 53.0 77.5 17.1 19.7 24.0 21.7 11.4 21.4 8.7 7.2
Biology/Genetics and Evolution 66.8 69.1 72.8 57.7 24.9 22.3 10.6 9.8 11.5 8.4 6.2 7.5
Biology/Physiology and Molecular Process 73.0 69.1 71.9 62.1 28.7 23.2 20.5 16.0 15.2 9.3 8.6 7.3
Biology/Structure and Morphology 76.5 74.2 74.1 72.4 46.8 39.0 33.6 28.6 27.5 19.1 16.2 21.7
Chemistry/Chemical Equilibrium 88.7 80.6 60.5 50.8 26.8 29.1 15.1 20.4 36.4 15.4 12.4 15.4
Chemistry/Chemical Kinetics and Thermochemistry 87.8 77.3 71.5 73.5 20.6 34.9 22.1 15.7 8.4 11.6 1.4 6.4
Chemistry/Chemical Reaction 34.2 33.7 26.4 30.2 10.4 3.5 4.9 3.9 5.0 2.3 0.3 2.0
Chemistry/Electrochemistry 74.4 78.3 79.8 70.4 20.1 24.1 16.8 10.1 15.3 12.3 2.2 11.3
Chemistry/Structure Of Matter 54.5 43.5 47.4 45.4 22.6 17.0 16.2 15.8 11.3 16.2 6.9 12.2
Computer/Data Structures and Algorithms 52.1 48.0 42.2 34.0 20.9 20.3 17.4 12.6 6.9 7.6 3.9 11.6
Computer/Hardware Architecture 48.2 49.4 48.2 41.7 14.8 20.1 12.2 5.8 7.9 2.4 3.0 2.4
Computer/Networking and Systems 59.8 57.0 50.4 50.6 19.1 10.9 12.9 9.5 3.1 1.9 2.7 0.9
Computer/Theory and AI 66.2 60.5 56.1 46.6 25.9 19.1 14.5 11.8 6.9 7.9 8.7 8.7
Economics/Finance and Decision 74.4 65.0 62.6 47.9 26.9 32.4 17.2 18.8 8.6 13.5 7.0 9.0
Economics/Macroeconomics 65.5 57.8 60.0 59.4 24.9 19.9 21.0 8.2 6.4 5.2 7.7 6.0
Economics/Microeconomics 64.8 52.0 59.0 62.1 18.0 18.9 15.9 8.8 10.1 6.7 6.1 11.0
Engineering/Civil and Architecture 66.4 62.4 60.5 64.7 33.3 34.9 30.2 18.2 21.8 13.9 16.0 13.8
Engineering/Mechanical Engineering 64.4 54.3 62.5 68.3 31.8 38.0 27.4 17.0 16.3 15.5 11.5 11.1
Engineering/Mechanics and Dynamics 61.7 51.1 48.3 45.6 11.1 15.3 9.7 0.0 2.3 0.6 8.5 2.6
Engineering/Mechanics of Materials 69.8 62.2 71.0 49.9 20.4 17.2 21.8 15.3 18.3 8.7 4.7 6.0
Engineering/Surveying and Cartography 74.5 65.4 79.5 72.0 37.1 37.8 33.7 11.6 16.9 12.8 20.9 15.3
Engineering/Thermodynamics and Energy 63.1 64.8 72.0 72.9 29.1 32.4 20.0 12.7 16.0 6.6 10.3 8.9
Geography/Earth Science 77.9 72.9 73.5 75.7 53.1 56.0 38.5 37.9 41.1 30.2 37.6 25.3
Geography/Human and Ecology 61.9 47.2 61.8 36.1 18.0 20.4 17.3 8.5 6.4 1.2 6.2 4.2
Geography/Maps 69.8 56.3 66.7 58.9 44.0 44.7 35.7 38.8 40.5 30.2 31.2 24.5
History/Historical Data Change 62.3 75.2 53.1 57.9 25.3 31.0 25.3 10.5 6.9 15.5 8.6 12.2
History/Historical Map 68.6 51.3 57.5 56.4 36.3 40.5 33.9 24.3 21.3 16.3 15.9 9.9
History/Others 65.0 79.1 64.8 78.8 18.3 34.8 19.5 19.8 5.3 12.8 13.1 7.8
Mathematics/Analytic Geometry 49.1 36.6 37.8 28.7 20.7 16.4 12.4 10.7 11.3 6.5 10.0 14.5
Mathematics/Plane Geometry 52.1 40.2 45.6 36.2 24.1 18.2 18.3 12.8 11.8 14.4 11.0 15.3
Mathematics/Solid Geometry 65.1 52.6 50.6 70.5 32.7 37.3 27.3 16.1 19.3 16.7 13.5 22.0
Music/Chord Structures and Interval Diagrams 36.6 31.2 32.6 21.2 12.2 8.1 13.5 9.1 6.3 1.0 1.0 5.9
Music/Key Signatures and Time Signatures 43.0 36.2 41.0 28.6 27.3 20.4 20.7 28.2 30.4 12.9 7.1 16.2
Music/Notes and Rests 57.7 34.9 37.2 43.7 27.9 27.2 23.1 22.8 26.8 17.9 10.7 16.9
Physics/Circuits and Electronics 55.1 36.8 47.3 39.1 14.5 14.7 8.8 9.7 5.0 7.2 4.6 3.7
Physics/Electromagnetism 71.9 60.0 66.0 66.0 29.3 35.1 26.0 19.1 19.1 24.6 18.2 20.9
Physics/Mechanics 74.9 47.4 79.3 71.5 33.1 37.7 23.6 11.4 16.8 14.0 16.2 15.6
Physics/Optics and Waves 74.5 66.2 63.0 69.8 38.3 28.8 24.8 36.0 21.5 20.7 19.0 7.9
Physics/Quantum Mechanics 63.4 60.1 58.8 64.8 35.7 28.6 25.1 20.7 26.9 24.1 19.5 13.3
Physics/Thermodynamics 68.8 50.6 53.2 51.6 17.7 22.9 8.2 14.0 3.4 5.1 7.5 7.8
Table 8: Relaxed scores on level-2 Taxonomy.
@@
Seedream 4.0
Gemini-2.5-Flash-Image
@@
BLIP3o-NEXT-GRPO-Text-3B
Biology Semantic Correctness 0.72 0.70 0.76 0.69 0.40 0.32 0.26 0.20 0.22 0.10 0.12 0.14
Spelling 1.53 1.50 0.93 1.21 0.49 0.28 0.25 0.29 0.22 0.17 0.04 0.13
Logical Consistency 1.32 1.13 1.24 1.17 0.50 0.51 0.31 0.43 0.40 0.39 0.40 0.29
Readability 1.74 1.71 1.72 1.63 1.15 1.15 1.24 1.07 0.69 1.01 0.37 0.67
Chemistry Semantic Correctness 0.46 0.41 0.43 0.43 0.15 0.14 0.10 0.07 0.08 0.08 0.04 0.06
Spelling 1.66 1.31 0.94 1.06 0.51 0.23 0.27 0.40 0.35 0.37 0.03 0.25
Logical Consistency 0.86 0.65 0.57 0.57 0.19 0.14 0.12 0.24 0.27 0.37 0.21 0.26
Readability 1.72 1.50 1.59 1.27 1.05 0.77 0.86 0.82 0.33 0.58 0.19 0.53
Economy Semantic Correctness 0.64 0.52 0.62 0.61 0.21 0.20 0.18 0.07 0.09 0.05 0.07 0.06
Spelling 1.66 1.71 1.36 1.34 0.42 0.39 0.19 0.29 0.08 0.00 0.01 0.17
Logical Consistency 1.06 0.64 0.65 0.78 0.19 0.16 0.19 0.08 0.06 0.19 0.29 0.29
Readability 1.53 1.52 1.32 1.27 0.95 0.69 0.88 0.48 0.21 0.36 0.14 0.27
Engineering Semantic Correctness 0.62 0.57 0.66 0.67 0.26 0.33 0.19 0.12 0.15 0.08 0.11 0.08
Spelling 1.64 1.60 1.30 1.36 0.67 0.50 0.66 0.32 0.43 0.23 0.12 0.24
Logical Consistency 1.17 0.89 1.05 1.01 0.42 0.41 0.35 0.21 0.27 0.34 0.41 0.18
Readability 1.64 1.49 1.55 1.41 1.02 0.92 1.15 0.60 0.46 0.48 0.31 0.44
Geography Semantic Correctness 0.72 0.64 0.71 0.68 0.48 0.49 0.30 0.29 0.34 0.20 0.32 0.21
Spelling 1.51 1.29 1.08 1.08 0.62 0.65 0.73 0.77 0.80 0.68 0.39 0.30
Logical Consistency 1.41 1.03 1.30 1.18 0.83 0.91 0.82 0.94 1.02 0.91 0.92 0.59
Readability 1.80 1.71 1.73 1.55 1.27 1.45 1.45 1.53 1.12 1.24 0.83 0.97
History Semantic Correctness 0.66 0.54 0.59 0.60 0.32 0.41 0.32 0.14 0.17 0.11 0.14 0.09
Spelling 1.41 1.32 0.93 0.83 0.40 0.39 0.27 0.39 0.27 0.15 0.12 0.02
Logical Consistency 1.15 0.85 0.78 0.80 0.55 0.39 0.39 0.71 0.34 0.66 0.49 0.15
Readability 1.63 1.68 1.49 1.46 1.20 1.24 1.20 1.34 0.68 0.90 0.34 0.56
Math Semantic Correctness 0.43 0.31 0.39 0.30 0.15 0.13 0.09 0.06 0.07 0.04 0.08 0.05
Spelling 1.77 1.60 1.41 1.38 1.12 0.81 0.83 0.54 0.79 0.53 0.18 0.90
Logical Consistency 0.94 0.54 0.48 0.40 0.36 0.19 0.27 0.26 0.26 0.58 0.52 0.62
Readability 1.62 1.49 1.21 1.16 1.08 0.90 0.93 0.77 0.44 0.59 0.38 0.81
Music Semantic Correctness 0.42 0.25 0.24 0.27 0.13 0.13 0.08 0.10 0.12 0.07 0.05 0.06
Spelling 1.74 1.42 1.46 1.51 1.38 0.91 0.88 0.78 1.18 0.86 0.57 0.63
Logical Consistency 0.98 0.45 0.77 0.74 0.51 0.42 0.75 0.71 1.09 0.23 0.18 0.52
Readability 1.89 1.49 1.74 1.63 1.42 1.49 1.54 1.40 0.80 0.86 0.28 1.09
Physics Semantic Correctness 0.62 0.43 0.61 0.57 0.22 0.25 0.12 0.12 0.11 0.09 0.12 0.07
Spelling 1.74 1.62 1.28 1.32 0.69 0.45 0.52 0.19 0.45 0.42 0.16 0.36
Logical Consistency 1.16 0.65 0.82 0.83 0.34 0.42 0.25 0.29 0.22 0.42 0.31 0.30
Readability 1.75 1.45 1.56 1.36 1.00 0.91 1.06 0.70 0.36 0.66 0.26 0.52
Computer
Science
Semantic Correctness 0.48 0.47 0.45 0.38 0.16 0.16 0.11 0.07 0.05 0.03 0.05 0.03
Spelling 1.75 1.73 1.35 1.25 0.87 0.66 0.52 0.32 0.35 0.29 0.05 0.51
Logical Consistency 1.05 0.71 0.58 0.47 0.23 0.21 0.25 0.27 0.10 0.24 0.14 0.28
Readability 1.66 1.44 1.30 1.04 0.84 0.75 0.81 0.58 0.18 0.29 0.11 0.45
@@ Appendix D Full Taxonomy List
In the table below, we show the complete list of four-level taxonomy and numbers of samples (denoted in parentheses). Blue text indicates mapping with ISCED-F UNESCO (2013) codes.
@@
Biology (156)
Ecology (5)
0521 Environmental sciences
Ecosystem (5) Food Chain and Food Web (4)
@@
Genetics and Evolution (13)
0511 Biology
Evolution and Population Genetics (3) Evolutionary Tree (2)
Pedigree Dominant and
Recessive Genetic Diseases (1)
@@
Transcription and
Splicing Mechanisms (1)
Physiology and Molecular Process (49)
0511 Biology
0512 Biochemistry
0912 Medicine
0916 Pharmacy
Cell Physiology (3) Cell Division (3)
@@
Structure and Morphology (89)
0511 Biology
0512 Biochemistry
Cell Structure (27) Basic Cell Structure (18)
@@
Chemistry (118)
0531 Chemistry
Chemical Equilibrium (5) Acid Base Titration (2)
@@
Chemical Kinetics
and Thermochemistry (7)
@@
Molecular Structure (49)
Electron Configuration and
Intermolecular Forces (8)
@@
Computer (102)
Data Structures and Algorithms (50)
0613 Software and
applications development and
analysis
ER Diagram (2)
@@
Hardware Architecture (17)
0613 Software and
applications development and
analysis
Bus Structure (3)
@@
Networking and Systems (11)
0612 Database and
network design and
administration
Computer Networks (7) Packet Structure (3)
@@
Operating System (4)
Deadlock Resource
Allocation Diagram (1)
@@
Theory and AI (24)
0613 Software and
applications development and
analysis
0619 Artificial Intelligence
Compiler Principles (2)
@@
Overfitting and
Underfitting Diagram (1)
@@
Economics (77)
Finance and Decision (5)
0412 Finance, banking
and insurance
CAPM (2) Beta Coefficient Scatter Plot (1)
@@
Macroeconomics (42)
0311 Economics
AD AS Model Aggregate
Demand and Aggregate Supply (29)
@@
Phillips Curve Inflation
and Unemployment (3)
Microeconomics (30)
0311 Economics
Cost Curves (4)
@@
Indifference Curve
Utility Maximization (6)
@@
Engineering (111)
Civil and Architecture (19)
0731 Architecture and
town planning
0732 Building and
civil engineering
Building Structure (4)
@@
Mechanical Engineering (33)
0715 Mechanics and
metal trades
Engineering Drawings (32) Schematic Diagram (18)
@@
Mechanics and Dynamics (9)
0715 Mechanics and
metal trades
Vibration and Control (9) Block Diagram (4)
@@
Mechanics of Materials (11)
0715 Mechanics and
metal trades
Stress Strain Relationship (11)
Surveying and Cartography (9)
0715 Mechanics and
metal trades
Data Processing and Adjustment (3)
@@
Thermodynamics and Energy (30)
0713 Electricity and
energy
Energy (2)
@@
Geography (66)
Earth Science (37)
0532 Earth sciences
Astronomy (3)
@@
Human and Ecology (4)
0314 Sociology and
cultural studies
Population and City (3)
@@
Maps (25)
0314 Sociology and
cultural studies
Earthquake Belt Distribution (2)
@@
History (41)
0222 History and
archaeology
Historical Data Change (7) Others (2)
@@
Mathematics (151)
0541 Mathematics
Analytic Geometry (56) Absolute Value Function (1)
@@
Exponential and
Logarithmic Function (2)
@@
Music (65)
0215 Music and
performing arts
Chord Structures and
Interval Diagrams (10)
Perfect Fifth and Major Third (3)
@@
Physics (113)
0533 Physics
0714 Electronics and
automation
Circuits and Electronics (36) Circuit Diagram (25) Curves In Circuit (2)
@@
Pressure Volume
Temperature Relationship (2)
Appendix E Limitations
While GenExam advances multidisciplinary T2I evaluation, its size (1000 samples) is still relatively small compared to multidisciplinary understanding benchmarks (e.g. 11.5K for MMMU) and may not cover all sub-disciplines and specified domains. In addition, the evaluation framework relies on MLLMs as judges – though enhanced by scoring points and ground truth, it still depends on the MLLMs’ own disciplinary knowledge and understanding capability, potentially introducing bias or errors in assessing niche domain details.
Appendix F Prompts
F.1 Automatic Filtering

### v3

@@ C.2 Detailed Comparison of Existing Methods
Biology/Ecology 100.0 86.9 47.1 26.1 47.8 77.5 40.6 19.7 24.0 21.4 8.7
Biology/Genetics_and_Evolution 94.9 87.6 66.8 75.9 64.4 57.7 36.2 22.3 10.6 8.4 6.2
Biology/Physiology_and_Molecular_Process 97.8 92.5 73.0 75.0 75.4 62.1 47.8 23.2 20.5 9.3 8.6
Biology/Structure_and_Morphology 94.7 92.5 76.5 79.0 77.0 72.4 62.2 39.0 33.6 19.1 16.2
Chemistry/Chemical_Equilibrium 100.0 97.3 88.7 83.7 67.0 50.8 56.3 29.1 15.1 15.4 12.4
Chemistry/Chemical_Kinetics_and_Thermochemistry 99.0 97.5 87.8 76.8 86.4 73.5 62.3 34.9 22.1 11.6 1.4
Chemistry/Chemical_Reaction 83.4 66.3 34.2 34.1 34.5 30.2 20.0 3.5 4.9 2.3 0.3
Chemistry/Electrochemistry 100.0 87.8 74.4 71.6 75.6 70.4 39.9 24.1 16.8 12.3 2.2
Chemistry/Structure_Of_Matter 88.4 79.5 54.5 48.5 56.9 45.4 34.1 17.0 16.2 16.2 6.9
Computer/Data_Structures_and_Algorithms 88.4 67.4 52.1 51.2 49.0 34.0 27.4 20.3 17.4 7.6 3.9
Computer/Hardware_Architecture 96.0 77.4 48.2 62.1 63.3 41.7 36.6 20.1 12.2 2.4 3.0
Computer/Networking_and_Systems 95.9 83.1 59.8 61.6 57.3 50.6 35.0 10.9 12.9 1.9 2.7
Computer/Theory_and_AI 93.7 88.9 66.2 67.0 66.8 46.6 32.9 19.1 14.5 7.9 8.7
Economics/Finance_and_Decision 92.0 86.8 74.4 68.1 72.6 47.9 37.6 32.4 17.2 13.5 7.0
Economics/Macroeconomics 97.7 83.5 65.5 67.4 61.3 59.4 43.9 19.9 21.0 5.2 7.7
Economics/Microeconomics 97.3 88.2 64.8 67.1 60.0 62.1 44.2 18.9 15.9 6.7 6.1
Engineering/Civil_and_Architecture 95.8 87.5 66.4 67.5 67.3 64.7 51.3 34.9 30.2 13.9 16.0
Engineering/Mechanical_Engineering 93.8 87.5 64.4 69.8 69.7 68.3 55.5 38.0 27.4 15.5 11.5
Engineering/Mechanics_and_Dynamics 95.6 84.0 61.7 70.5 40.2 45.6 44.9 15.3 9.7 0.6 8.5
Engineering/Mechanics_of_Materials 96.1 89.5 69.8 68.4 68.4 49.9 33.8 17.2 21.8 8.7 4.7
Engineering/Surveying_and_Cartography 98.9 84.6 74.5 68.7 80.0 72.0 60.1 37.8 33.7 12.8 20.9
Engineering/Thermodynamics_and_Energy 94.5 84.7 63.1 71.6 74.4 72.9 43.4 32.4 20.0 6.6 10.3
Geography/Earth_Science 97.8 94.5 77.9 75.8 82.8 75.7 68.7 56.0 38.5 30.2 37.6
Geography/Human_and_Ecology 93.5 98.2 61.9 51.7 50.2 36.1 22.9 20.4 17.3 1.2 6.2
Geography/Maps 94.9 88.7 69.8 58.1 70.8 58.9 60.0 44.7 35.7 30.2 31.2
History/Historical_Data_Change 100.0 85.3 62.3 71.8 77.7 57.9 48.6 31.0 25.3 15.5 8.6
History/Historical_Map 99.8 94.8 68.6 50.1 64.5 56.4 47.4 40.5 33.9 16.3 15.9
History/Others 100.0 49.2 65.0 74.8 91.1 78.8 45.2 34.8 19.5 12.8 13.1
Mathematics/Analytic_Geometry 90.7 68.1 49.1 45.7 38.7 28.7 25.3 16.4 12.4 6.5 10.0
Mathematics/Plane_Geometry 83.7 63.0 52.1 43.4 54.0 36.2 33.8 18.2 18.3 14.4 11.0
Mathematics/Solid_Geometry 83.9 76.0 65.1 50.0 63.6 70.5 47.1 37.3 27.3 16.7 13.5
Music/Chord_Structures_and_Interval_Diagrams 93.4 52.4 36.6 48.1 36.4 21.2 19.1 8.1 13.5 1.0 1.0
Music/Key_Signatures_and_Time_Signatures 85.4 55.7 43.0 29.7 37.5 28.6 28.3 20.4 20.7 12.9 7.1
Music/Notes_and_Rests 91.5 77.2 57.7 37.3 50.9 43.7 37.3 27.2 23.1 17.9 10.7
Physics/Circuits_and_Electronics 90.2 73.2 55.1 55.6 47.1 39.1 29.1 14.7 8.8 7.2 4.6
Physics/Electromagnetism 97.5 91.9 71.9 66.1 77.4 66.0 53.0 35.1 26.0 24.6 18.2
Physics/Mechanics 99.4 92.9 74.9 68.2 76.7 71.5 50.7 37.7 23.6 14.0 16.2
Physics/Optics_and_Waves 95.2 87.7 74.5 74.9 71.1 69.8 54.7 28.8 24.8 20.7 19.0
Physics/Quantum_Mechanics 94.9 85.1 63.4 63.5 66.4 64.8 46.1 28.6 25.1 24.1 19.5
Physics/Thermodynamics 96.6 93.5 68.8 61.9 53.4 51.6 38.9 22.9 8.2 5.1 7.5
Table 8: Relaxed scores on level-2 raxonomy.
@@
Seedream 4.5
FLUX.2 max
@@
Biology Semantic Correctness 0.96 0.91 0.72 0.74 0.78 0.69 0.59 0.32 0.26 0.1 0.12
Spelling 1.9 1.95 1.53 1.72 1.05 1.21 0.7 0.28 0.25 0.17 0.04
Logical Consistency 1.91 1.76 1.32 1.32 1.3 1.17 0.73 0.51 0.31 0.39 0.4
Readability 1.98 1.94 1.74 1.74 1.69 1.63 1.28 1.15 1.24 1.01 0.37
Chemistry Semantic Correctness 0.86 0.75 0.46 0.42 0.53 0.43 0.3 0.14 0.1 0.08 0.04
Spelling 1.97 1.87 1.66 1.59 1.21 1.06 0.88 0.23 0.27 0.37 0.03
Logical Consistency 1.72 1.37 0.86 0.69 0.72 0.57 0.46 0.14 0.12 0.37 0.21
Readability 1.96 1.94 1.72 1.6 1.43 1.27 1.15 0.77 0.86 0.58 0.19
Economics Semantic Correctness 0.97 0.84 0.64 0.65 0.64 0.61 0.44 0.2 0.18 0.05 0.07
Spelling 2.0 1.94 1.66 1.88 1.35 1.34 1.06 0.39 0.19 0.0 0.01
Logical Consistency 1.95 1.52 1.06 0.88 0.69 0.78 0.44 0.16 0.19 0.19 0.29
Readability 1.97 1.82 1.53 1.57 1.29 1.27 1.1 0.69 0.88 0.36 0.14
Engineering Semantic Correctness 0.95 0.84 0.62 0.68 0.71 0.67 0.5 0.33 0.19 0.08 0.11
Spelling 1.95 1.97 1.64 1.79 1.29 1.36 1.0 0.5 0.66 0.23 0.12
Logical Consistency 1.85 1.64 1.17 1.08 1.05 1.01 0.59 0.41 0.35 0.34 0.41
Readability 1.95 1.86 1.64 1.56 1.46 1.41 1.19 0.92 1.15 0.48 0.31
Geography Semantic Correctness 0.96 0.9 0.72 0.65 0.78 0.68 0.66 0.49 0.3 0.2 0.32
Spelling 1.95 1.97 1.51 1.61 1.3 1.08 0.82 0.65 0.73 0.68 0.39
Logical Consistency 1.89 1.88 1.41 1.12 1.32 1.18 1.03 0.91 0.82 0.91 0.92
Readability 1.95 2.0 1.8 1.65 1.7 1.55 1.47 1.45 1.45 1.24 0.83
History Semantic Correctness 1.0 0.89 0.66 0.51 0.71 0.6 0.52 0.41 0.32 0.11 0.14
Spelling 2.0 2.0 1.41 1.56 1.1 0.83 0.56 0.39 0.27 0.15 0.12
Logical Consistency 2.0 1.76 1.15 0.83 0.98 0.8 0.56 0.39 0.39 0.66 0.49
Readability 2.0 1.95 1.63 1.49 1.66 1.46 1.12 1.24 1.2 0.9 0.34
Math Semantic Correctness 0.84 0.6 0.43 0.37 0.46 0.3 0.26 0.13 0.09 0.04 0.08
Spelling 1.97 1.95 1.77 1.73 1.53 1.38 1.28 0.81 0.83 0.53 0.18
Logical Consistency 1.66 1.12 0.94 0.59 0.61 0.4 0.3 0.19 0.27 0.58 0.52
Readability 1.89 1.7 1.62 1.45 1.3 1.16 1.08 0.9 0.93 0.59 0.38
Music Semantic Correctness 0.9 0.64 0.42 0.28 0.36 0.27 0.23 0.13 0.08 0.07 0.05
Spelling 1.98 1.97 1.74 1.62 1.63 1.51 1.38 0.91 0.88 0.86 0.57
Logical Consistency 1.74 1.26 0.98 0.54 1.02 0.74 0.57 0.42 0.75 0.23 0.18
Readability 1.92 1.89 1.89 1.57 1.71 1.63 1.54 1.49 1.54 0.86 0.28
Physics Semantic Correctness 0.95 0.83 0.62 0.6 0.63 0.57 0.42 0.25 0.12 0.09 0.12
Spelling 1.97 1.96 1.74 1.82 1.42 1.32 1.03 0.45 0.52 0.42 0.16
Logical Consistency 1.84 1.65 1.16 0.85 0.88 0.83 0.5 0.42 0.25 0.42 0.31
Readability 1.96 1.9 1.75 1.64 1.52 1.36 1.12 0.91 1.06 0.66 0.26
Computer Science Semantic Correctness 0.91 0.72 0.48 0.53 0.54 0.38 0.27 0.16 0.11 0.03 0.05
Spelling 1.98 1.94 1.75 1.84 1.49 1.25 1.1 0.66 0.52 0.29 0.05
Logical Consistency 1.75 1.31 1.05 0.8 0.84 0.47 0.33 0.21 0.25 0.24 0.14
Readability 1.93 1.84 1.66 1.51 1.42 1.04 0.97 0.75 0.81 0.29 0.11
@@ Appendix D Full Taxonomy List
In the table below, we show the complete list of four-level taxonomy and numbers of samples (denoted in parentheses). Blue text indicates mapping with ISCED-F (UNESCO, 2013) codes.
@@
Biology (156) Ecology (5) 0521 Environmental sciences Ecosystem (5) Food Chain and Food Web (4)
@@
Genetics and Evolution (13) 0511 Biology Evolution and Population Genetics (3) Evolutionary Tree (2)
Pedigree Dominant and Recessive Genetic Diseases (1)
@@
Transcription and Splicing Mechanisms (1)
Physiology and Molecular Process (49) 0511 Biology 0512 Biochemistry 0912 Medicine 0916 Pharmacy Cell Physiology (3) Cell Division (3)
@@
Structure and Morphology (89) 0511 Biology 0512 Biochemistry Cell Structure (27) Basic Cell Structure (18)
@@
Chemistry (118) 0531 Chemistry Chemical Equilibrium (5) Acid Base Titration (2)
@@
Chemical Kinetics and Thermochemistry (7)
@@
Molecular Structure (49) Electron Configuration and Intermolecular Forces (8)
@@
Computer (102) Data Structures and Algorithms (50) 0613 Software and applications development and analysis ER Diagram (2)
@@
Hardware Architecture (17) 0613 Software and applications development and analysis Bus Structure (3)
@@
Networking and Systems (11) 0612 Database and network design and administration Computer Networks (7) Packet Structure (3)
@@
Operating System (4) Deadlock Resource Allocation Diagram (1)
@@
Theory and AI (24) 0613 Software and applications development and analysis 0619 Artificial Intelligence Compiler Principles (2)
@@
Overfitting and Underfitting Diagram (1)
@@
Economics (77) Finance and Decision (5) 0412 Finance, banking and insurance CAPM (2) Beta Coefficient Scatter Plot (1)
@@
Macroeconomics (42) 0311 Economics AD AS Model Aggregate Demand and Aggregate Supply (29)
@@
Phillips Curve Inflation and Unemployment (3)
Microeconomics (30) 0311 Economics Cost Curves (4)
@@
Indifference Curve Utility Maximization (6)
@@
Engineering (111) Civil and Architecture (19) 0731 Architecture and town planning 0732 Building and civil engineering Building Structure (4)
@@
Mechanical Engineering (33) 0715 Mechanics and metal trades Engineering Drawings (32) Schematic Diagram (18)
@@
Mechanics and Dynamics (9) 0715 Mechanics and metal trades Vibration and Control (9) Block Diagram (4)
@@
Mechanics of Materials (11) 0715 Mechanics and metal trades Stress Strain Relationship (11)
Surveying and Cartography (9) 0715 Mechanics and metal trades Data Processing and Adjustment (3)
@@
Thermodynamics and Energy (30) 0713 Electricity and energy Energy (2)
@@
Geography (66) Earth Science (37) 0532 Earth sciences Astronomy (3)
@@
Human and Ecology (4) 0314 Sociology and cultural studies Population and City (3)
@@
Maps (25) 0314 Sociology and cultural studies Earthquake Belt Distribution (2)
@@
History (41) 0222 History and archaeology Historical Data Change (7) Others (2)
@@
Mathematics (151) 0541 Mathematics Analytic Geometry (56) Absolute Value Function (1)
@@
Exponential and Logarithmic Function (2)
@@
Music (65) 0215 Music and performing arts Chord Structures and Interval Diagrams (10) Perfect Fifth and Major Third (3)
@@
Physics (113) 0533 Physics 0714 Electronics and automation Circuits and Electronics (36) Circuit Diagram (25) Curves In Circuit (2)
@@
Pressure Volume Temperature Relationship (2)
Appendix E Prompts
E.1 Automatic Filtering